# Cross-Domain Few-Shot Learning for Hyperspectral Image Classification Based on Mixup Foundation Model

### v1

@@ III-B Mixup Strategy
III-C Prototypical Network
MIFOMO is built upon the idea of prototypical network [44] using the prototypes to classify the query samples. The prototypes are the centers of the samples of each class, quantified as follows:
@@ III-C Prototypical Network
IV Methodology
MIFOMO is constructed under a hyper-spectral foundation model, namely HyperSIGMA [47], pretrained using the Hyper-Global-450K dataset by means of the masked image modeling technique. HyperSIGMA is a vision transformer (ViT)-based foundation model [9], and its innovation lies in the proposal of the sparse sampling attention (SSA) method instead of the self-attention mechanism. Unlike [3], using the natural-images foundation model, HyperSIGMA is meant specifically to handle HSIs. We propose a coalescent projection (CP) method as a parameter-efficient fine-tuning (PEFT) technique while leaving the HyperSIGMA backbone completely frozen. The CP method establishes a single learnable matrix that embraces the query-key interactions. We put forward the mixup strategy where the episodic-based meta-learning phase occurs in the source and intermediate domains. The intermediate domain is crafted from the mixup samples between the source and target domains and functions as a bridge between the two domains for the sake of seamless knowledge transfer.
IV-A Network Structure of HyperSIGMA
HyperSIGMA [47] comprises two sub-networks: SpatialNetwork and SpectralNetwork, where spatial and spectral features are ultimately fused to enhance the representation of the extracted features. Fig. 2 visualizes the architecture of MIFOMO. Both networks are based on the ViT backbone. However, the spatial tokenization is extended to the spectral domain, generating spectral tokens by embedding channels. That is, a 3-D HSI cube $x_{0}\in\Re^{H\times W\times C}$ is aggregated through the average clustering technique along the channel dimension, leading to $X^{{}^{\prime}}\in\Re^{H\times W\times N_{\text{spec}}}$ where $N_{\text{spec}}$ is the desired token number. The dimensional permutation and flattening are applied to reshape $X^{{}^{\prime}}$ into a 2-D matrix of $\Re^{N_{\text{spec}}\times(H.W)}$. It is then embedded through a linear mapping and generates $X\in^{N_{\text{spec}}\times D}$, being akin to the spatial patch embedding in normal ViT.
@@ IV-A Network Structure of HyperSIGMA
where $h$ is the number of heads and $W\in\Re^{hD^{{}^{\prime}}\times D}$ is a linear projector to recover the original embedding dimension $D$. We choose to retain the SA mechanism for all layers rather than using the SSA method because the SSA method is not applied in all layers in HyperSIGMA. It includes replacing the original SA with SSA, complicating the fine-tuning process.
HyperSIGMA possesses a dual-branch network structure that produces spectral and spatial features, respectively. The two features are fused to enhance their representations. To this end, the spectral enhancement module (SEM) is designed to enhance spatial features with spectral information. Suppose that we have a spatial feature $Z_{\text{spat}}\in\Re^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times D}$ and a spectral feature $Z_{\text{spec}}\in\Re^{N_{\text{spec}}\times D}$, a linear layer is applied to reduce their dimensions to $D_{1}$. We further reduce the spatial dimension of $Z_{\text{spec}}$ via average aggregation to produce a $1D$ vector $Z^{{}^{\prime}}\in\Re^{N_{\text{spec}}}$. Lastly, another linear mapping is implemented to align the dimension of $Z^{{}^{\prime}}$ with the channel number of $Z_{\text{spat}}$. The spectral enhancement process is formalised:
@@
IV-B Coalescent Projection
MIFOMO is equipped with a coalescent projection (CP) seen as a generalization of the soft prompt [52]. We start our discussion by introducing the attention map, being a crucial part of the self-attention mechanism as per (7). It can be decomposed into 4 blocks.
@@ IV-B Coalescent Projection
From [1], we understand that the most crucial part of the attention map calculation is $A_{1,1}$ because the prompt can only affect $A_{1,1}$ and $A_{1,p}$ due to the presence of the softmax function. Since the prompt is injected in every layer to correct the influence of the prompt in the previous layer, the bottom part of the attention map $A$ is deemed redundant. Figure 1 exhibits the self-attention mechanism of the plain prompt [52] and the CP in this paper.
@@
Compared to the prompt method [52] inserting external trainable parameters to the patch embedding, the CP concent has at least three clear advantages: 1) the CP method doesn’t have any specific hyper-parameter whereas the prompt method one has to select an appropriate length of the prompt token; 2) the prompts are shared across all heads hampering model’s generalization due to the gradient’s interference whereas the CP concept handles each head independently; 3) the prompt increases the size of patch dimension consuming the context window and memory whereas the CP technique doesn’t manipulate the calculation size from that of the original attention calculation; 4) the CP method can change the order of the original input tokens whereas the prompt doesn’t have such aptitude. On the other hand, the CP method also imposes significantly fewer parameters than the attention scale [1]. This is especially important for high-resolution images such as HSIs. In addition, the attention scale method is not applicable to different input tokens.
@@
IV-C Mixup Domain Adaptation Strategy
Motivated by [11], the mixup domain adaptation strategy is proposed to address the domain shift problem in the CDFSL for the HSI classification problem. This strategy consists of two phases, namely the source domain and the intermediate domain. Note that the episodic-based meta-learning strategy is implemented rather than the conventional training paradigm.
IV-C1 Source Domain
@@
(15) is solved using the episodic-based meta-learning strategy for both original and mixup samples. That is, the loss is calculated from the query set of both original and mixup samples. This strategy enhances the robustness of the network because it aids the network in reserving the uncharted latent space uncovered by the original samples [34]. That is, it increases the complexity of the decision boundaries. Fig. 3 shows the flowchart of the source domain training phase.
IV-C2 Intermediate Domain
@@
The first step is performed by assigning high-quality pseudo-labels to the target domain’s query set. Next, the network is trained on the intermediate domain.
To assign pseudo-labels, the network is trained only on the support set of the target domain. This is done by sampling smaller support set $\mathcal{\overline{S}}_{S^{T}}=\{(x_{i},y_{i})\}_{i=1}^{N\times K_{s}}$ and query sets $\mathcal{\overline{Q}}_{S_{T}}=\{(x_{i},y_{i})\}_{i=1}^{N\times K_{q}}$ per episode as subsets of the current support set and using mixup in the feature space, as in the source-domain phase, where $K_{s}$ and $K_{q}$ are the number of support and query subset shots, respectively, and $K_{s}+K_{q}=K$.
@@
The pseudo-label is generated based on a small number of samples of the support set of the target domain. That is, the prototypes are calculated based on the support set of the target domain $\mathcal{S}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N\times K}$. However, these pseudo-labels are noisy because of the domain shift problem, i.e., the network is not yet trained to the target domain, and the target domain shares no overlapping label space with the source domain. To this end, the label smoothing strategy [69] is utilized and enforces consistent predictions of adjacent samples. That is, neighboring samples must be labeled similarly or be smooth. A graph $G=(V,E)$ is constructed where $E$ is the nodes of the graph defined by all images in $\mathcal{Q}_{t}$ and $E$ is the edges weighted by an adjacency matrix $A\in\Re^{N_{Q}\times N_{Q}}$ whose elements are derived from the distance between two images in the feature space $a_{i,j}=\exp{(-\frac{(g_{\psi}(x_{i})-g_{\psi}(x_{j}))^{2}}{2\sigma^{2}})},j\neq i,a_{i,i}=0$. $\sigma$ is the spread of the latent space. The adjacency matrix is normalized as per $\hat{A}=D^{-1/2}AD^{-1/2}$ in which $D$ is a diagonal matrix whose diagonal element stands for the sum of the ithi-th row of $A$. The label propagation is iterated until convergence $\mathcal{F}^{*}$.
@@
The mixup ratio $\lambda_{2}$ plays a vital role in steering the relationship of the mixup samples to the source and target domains, determining knowledge transfer. We follow the progressive mixup strategy [71], where the Wasserstein distance $d(.)$ is applied to measure the discrepancies between the two domains, i.e., intermediate to source and intermediate to target. Initially, the intermediate domain is set to be close to the target domain. The mixup ratio is gradually adjusted to bring the intermediate domain closer to the source domain, thereby achieving domain alignment. This setting reduces large domain gaps because the network is trained to master the target domain and subsequently the source domain, thus ensuring seamless knowledge transfer. This is attained by introducing the weighting factor $q$ to portray the similarity between the two domains.
@@ IV-D Algorithm
IV-E Time-complexity Analysis
@@
V Experiments
V-A Datasets Specifications
To evaluate the effectiveness of the proposed method, we use a widely used, publicly available CDFSL for HSI classification datasets. We use the Chikusei dataset (CH) [59] for the source domain. Moreover, we use the Indian Pines (IP), Salinas (SA), and Pavia University (PU) datasets for the target domain [24].
The CH dataset was collected using the Headwall Hyperpec-VNIR-C sensor in Chikusei, Ibaraki, Japan, on July 29, 2014, which comprises 128 bands spanning 340-1018 nm. The image is 2517x2335 and contains 19 categories. More details are provided in table I.
The IP dataset was acquired by AVIRIS in 1992 over Northwest Indiana and contains 200 bands spanning 400-2500 nm, with a spatial size of 145x145 and 16 classes. The SA dataset was taken by the AVIRIS sensor in the Salinas Valley, California, USA. The image size is 512x217. The spectral image contains 224 bands, of which 204 bands are useful. The PU dataset was taken by the ROSIS sensor over Pavia, in northern Italy, in 2003. It contains 103 bands after discarding the noise bands, and covers the range 430-860 nm. The image size is 620x340 and has 9 classes. More details about the target datasets are provided in Tables I-V [42, 25].
TABLE I: Class names and samples per class for Chikusei dataset
@@ 19 Paved ground 145
TABLE II: Class names and samples per class for the Indian Pines dataset
@@ 16 Stone Steel Towers 93
TABLE III: Class names and samples per class for the Salinas dataset
@@ 16 Vinyard untrained trellis 1807
TABLE IV: Class names and samples per class for the Pavia University dataset
@@ 9 Shadows 947
TABLE V: Number of training and testing set samples per class in the Houston dataset
@@ 15 Running Track 187 473
V-B Baseline methods
To evaluate the effectiveness of our method, we comprehensively compare it against the MGPDO [42], CFSL-KT [17], CDLA [60], CF-Trans [70], CMTL [5], GMTN [49], DCFSL [24], CMFSL [54], HFSL [50], CNN-IT [16], Gia-CFSL [65], SSRN [68], MCNN [67], SPRN [64], FDFSL [38], ADAFSL [57], RPCL-FSL [28], CASCL [25], GLGAT-CFSL [7], and DA-CFSL [39].
V-C Implementation Details
The experiments are conducted on a workstation with an NVIDIA GeForce RTX 4090 GPU and Intel Core i9-14900K. We performed the experiments with 20 episodes (random seeds). We utilize the ViT-B model from HyperSIGMA as the backbone with two ViT branches pre-trained with masked autoencoder (MAE) [13] on the HyperGlobal-450K dataset [47]. Moreover, we use Principal Component Analysis (PCA) to decrease the number of bands to 50 for all datasets as a preprocessing step.
TABLE VI: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Indian pines dataset (Five labeled samples per class)
@@ V-C Implementation Details
SSRN [68] 99.01 42.91 58.84 81.49 78.63 95.75 99.55 76.91 100 60.51 51.14 53.59 99.29 87.98 75.21 97.71 $65.72\pm 3.77$ $79.22\pm 1.62$ $61.72\pm 3.90$
MCNN [67] 93.24 46.14 46.24 76.57 66.39 87.15 97.27 78.14 100 65.32 62.34 53.17 97.14 87.24 75.96 97.58 $66.17\pm 3.21$ $80.32\pm 2.83$ $62.57\pm 3.67$
SPRN [64] 98.94 54.76 51.14 83.78 73.53 90.91 99.65 84.13 100 67.95 65.22 49.22 97.58 90.22 80.18 96.51 $67.32\pm 2.08$ $81.28\pm 1.12$ $63.32\pm 2.93$
DCFSL [24] 97.8 45.14 46.88 75.60 67.74 86.50 99.57 76.03 100 57.08 61.45 48.69 95.35 89.48 67.72 96.82 $64.89\pm 3.50$ $75.75\pm 2.44$ $60.36\pm 3.96$
Gia-CFSL [65] 93.41 47.98 49.35 76.51 66.65 84.14 99.57 83.76 100 58.23 61.32 45.83 96.45 87.83 70.21 96.02 $65.44\pm 2.23$ $76.08\pm 1.48$ $61.02\pm 2.53$
FDFSL [38] 96.59 55.62 64.05 97.07 75.17 83.52 99.57 83.11 100 66.10 64.51 59.29 98.70 88.60 80.21 98.18 $71.11\pm 2.19$ $81.27\pm 1.84$ $67.55\pm 2.44$
ADAFSL [57] 100 51.03 64.10 81.21 79.75 92.65 100 92.30 100 59.54 69.93 59.93 97.20 87.86 71.50 97.73 $71.91\pm 5.15$ $81.54\pm 1.49$ $68.23\pm 3.39$
RPCL-FSL [28] 98.54 62.58 57.09 93.23 78.16 90.06 100 90.51 100 73.80 69.56 64.40 98.50 90.55 80.63 97.05 $75.11\pm 1.97$ $84.04\pm 1.10$ $71.94\pm 2.17$
CFSL-KT [17] 98.57 44.86 72.30 88.45 74.94 96.80 100.0 99.24 100.0 78.82 73.99 61.87 100.0 91.71 83.66 99.77 $76.29\pm 1.26$ $85.36\pm 1.21$ $73.41\pm 1.39$
GLGAT-CFSL [7] 100 78.83 91.93 95.79 96.99 95.03 99.74 98.91 99.67 79.82 80.96 70.03 98.75 97.84 68.93 100 78.5 82.96 79.98
CDLA [60] 99.27 68.01 73.56 92.24 80.77 93.23 100.0 95.18 100.0 71.61 70.99 71.43 97.90 92.49 91.39 98.07 $78.94\pm 3.88$ $87.26\pm 1.59$ $76.26\pm 4.23$
DA-CFSL [39] 99.27 73.23 73.78 91.25 82.34 91.75 100 94.29 100 71.32 70.87 74.78 97.9 90.71 84.78 98.75 $79.26\pm 3.45$ $87.19\pm 1.29$ $76.62\pm 3.73$
CASCL [25] 97.56 73.11 62.85 95.32 77.86 94.86 100 95.32 100 76.25 78.14 63.20 98.33 89.13 86.75 97.78 $79.73\pm 1.48$ $86.34\pm 1.73$ $76.53\pm 1.86$
MGPDO [42] 100 71.98 63.56 90.37 83.47 96.37 100 99.93 100 75.32 80.98 59.41 99.33 92.67 91.69 95.45 $81.32\pm 1.62$ $87.53\pm 0.31$ $78.75\pm 1.73$
@@
TABLE VII: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Pavia University dataset (Five labeled samples per class)

### v2

@@ 3.2 Mixup Strategy
3.3 Prototypical Network
MIFOMO is built upon the idea of prototypical network Snell2017PrototypicalNF using the prototypes to classify the query samples. The prototypes are the centers of the samples of each class, quantified as follows:
@@ 3.3 Prototypical Network
4 Methodology
MIFOMO is constructed under a hyper-spectral foundation model, namely HyperSIGMA Wang2024HyperSIGMAHI, pretrained using the Hyper-Global-450K dataset by means of the masked image modeling technique. HyperSIGMA is a vision transformer (ViT)-based foundation model Dosovitskiy2020AnII, and its innovation lies in the proposal of the sparse sampling attention (SSA) method instead of the self-attention mechanism. Unlike Chen2025SpectralDINODM, using the natural-images foundation model, HyperSIGMA is meant specifically to handle HSIs. We propose a coalescent projection (CP) method as a parameter-efficient fine-tuning (PEFT) technique while leaving the HyperSIGMA backbone completely frozen. The CP method establishes a single learnable matrix that embraces the query-key interactions. We put forward the mixup strategy where the episodic-based meta-learning phase occurs in the source and intermediate domains. The intermediate domain is crafted from the mixup samples between the source and target domains and functions as a bridge between the two domains for the sake of seamless knowledge transfer.
4.1 Network Structure of HyperSIGMA
HyperSIGMA Wang2024HyperSIGMAHI comprises two sub-networks: SpatialNetwork and SpectralNetwork, where spatial and spectral features are ultimately fused to enhance the representation of the extracted features. Fig. 2 visualizes the architecture of MIFOMO. Both networks are based on the ViT backbone. However, the spatial tokenization is extended to the spectral domain, generating spectral tokens by embedding channels. That is, a 3-D HSI cube $x_{0}\in\Re^{H\times W\times C}$ is aggregated through the average clustering technique along the channel dimension, leading to $X^{{}^{\prime}}\in\Re^{H\times W\times N_{\text{spec}}}$ where $N_{\text{spec}}$ is the desired token number. The dimensional permutation and flattening are applied to reshape $X^{{}^{\prime}}$ into a 2-D matrix of $\Re^{N_{\text{spec}}\times(H.W)}$. It is then embedded through a linear mapping and generates $X\in^{N_{\text{spec}}\times D}$, being akin to the spatial patch embedding in normal ViT.
@@ 4.1 Network Structure of HyperSIGMA
where $h$ is the number of heads and $W\in\Re^{hD^{{}^{\prime}}\times D}$ is a linear projector to recover the original embedding dimension $D$. We choose to retain the SA mechanism for all layers rather than using the SSA method because the SSA method is not applied to all layers in HyperSIGMA. It includes replacing the original SA with SSA, complicating the fine-tuning process.
HyperSIGMA possesses a dual-branch network structure that produces spectral and spatial features, respectively. The two features are fused to enhance their representations. To this end, the spectral enhancement module (SEM) is designed to enhance spatial features with spectral information. Suppose that we have a spatial feature $Z_{\text{spat}}\in\Re^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times D}$ and a spectral feature $Z_{\text{spec}}\in\Re^{N_{\text{spec}}\times D}$, a linear layer is applied to reduce their dimensions to $D_{1}$. We further reduce the spatial dimension of $Z_{\text{spec}}$ via average aggregation to produce a $1D$ vector $Z^{{}^{\prime}}\in\Re^{N_{\text{spec}}}$. Lastly, another linear mapping is implemented to align the dimension of $Z^{{}^{\prime}}$ with the channel number of $Z_{\text{spat}}$. The spectral enhancement process is formalized:
@@
4.2 Coalescent Projection
MIFOMO is equipped with a coalescent projection (CP) seen as a generalization of the soft prompt Wang2021LearningTP. We start our discussion by introducing the attention map, being a crucial part of the self-attention mechanism as per (7). It can be decomposed into 4 blocks.
@@ 4.2 Coalescent Projection
From Basu2023StrongBF, we understand that the most crucial part of the attention map calculation is $A_{1,1}$ because the prompt can only affect $A_{1,1}$ and $A_{1,p}$ due to the presence of the softmax function. Since the prompt is injected in every layer to correct the influence of the prompt in the previous layer, the bottom part of the attention map $A$ is deemed redundant. Figure 1 exhibits the self-attention mechanism of the plain prompt Wang2021LearningTP and the CP in this paper.
@@
Compared to the prompt method Wang2021LearningTP inserting external trainable parameters to the patch embedding, the CP concent has at least three clear advantages: 1) the CP method doesn’t have any specific hyper-parameter whereas in the prompt method one has to select an appropriate length of the prompt token; 2) the prompts are shared across all heads hampering model’s generalization due to the gradient’s interference whereas the CP concept handles each head independently; 3) the prompt increases the size of patch dimension consuming the context window and memory whereas the CP technique doesn’t manipulate the calculation size from that of the original attention calculation; 4) the CP method can change the order of the original input tokens whereas the prompt doesn’t have such aptitude. On the other hand, the CP method also imposes significantly fewer parameters than the attention scale Basu2023StrongBF. This is especially important for high-resolution images such as HSIs. In addition, the attention scale method is not applicable to different input tokens.
@@
4.3 Mixup Domain Adaptation Strategy
Motivated by Furqon2024MixupDA, the mixup domain adaptation strategy is proposed to address the domain shift problem in the CDFSL for the HSI classification problem. This strategy consists of two phases, namely the source domain and the intermediate domain. Note that the episodic-based meta-learning strategy is implemented rather than the conventional training paradigm.
4.3.1 Source Domain
@@
(15) is solved using the episodic-based meta-learning strategy for both original and mixup samples. That is, the loss is calculated from the query set of both original and mixup samples. This strategy enhances the robustness of the network because it aids the network in reserving the uncharted latent space uncovered by the original samples Paeedeh2025CrossDomainFL. That is, it increases the complexity of the decision boundaries. Fig. 3 shows the flowchart of the source domain training phase.
4.3.2 Intermediate Domain
@@
The first step is performed by assigning high-quality pseudo-labels to the target domain’s query set. Next, the network is trained on the intermediate domain. To assign pseudo-labels, the network is trained only on the support set of the target domain. This is done by sampling smaller support set $\mathcal{\overline{S}}_{S^{T}}=\{(x_{i},y_{i})\}_{i=1}^{N\times K_{s}}$ and query sets $\mathcal{\overline{Q}}_{S_{T}}=\{(x_{i},y_{i})\}_{i=1}^{N\times K_{q}}$ per episode as subsets of the current support set and using mixup in the feature space, as in the source-domain phase, where $K_{s}$ and $K_{q}$ are the number of support and query subset shots, respectively, and $K_{s}+K_{q}=K$.
@@
The pseudo-label is generated based on a small number of samples of the support set of the target domain. That is, the prototypes are calculated based on the support set of the target domain $\mathcal{S}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N\times K}$. However, these pseudo-labels are noisy because of the domain shift problem, i.e., the network is not yet trained to the target domain, and the target domain shares no overlapping label space with the source domain. To this end, the label smoothing strategy Zhou2003LearningWL is utilized and enforces consistent predictions of adjacent samples. That is, neighboring samples must be labeled similarly or be smooth. A graph $G=(V,E)$ is constructed where $E$ is the nodes of the graph defined by all images in $\mathcal{Q}_{t}$ and $E$ is the edges weighted by an adjacency matrix $A\in\Re^{N_{Q}\times N_{Q}}$ whose elements are derived from the distance between two images in the feature space $a_{i,j}=\exp{(-\frac{(g_{\psi}(x_{i})-g_{\psi}(x_{j}))^{2}}{2\sigma^{2}})},j\neq i,a_{i,i}=0$. $\sigma$ is the spread of the latent space. The adjacency matrix is normalized as per $\hat{A}=D^{-1/2}AD^{-1/2}$ in which $D$ is a diagonal matrix whose diagonal element stands for the sum of the ithi-th row of $A$. The label propagation is iterated until convergence $\mathcal{F}^{*}$.
@@
The mixup ratio $\lambda_{2}$ plays a vital role in steering the relationship of the mixup samples to the source and target domains, determining knowledge transfer. We follow the progressive mixup strategy Zhu2023ProgressiveMF, where the Wasserstein distance $d(.)$ is applied to measure the discrepancies between the two domains, i.e., intermediate to source and intermediate to target. Initially, the intermediate domain is set to be close to the target domain. The mixup ratio is gradually adjusted to bring the intermediate domain closer to the source domain, thereby achieving domain alignment. This setting reduces large domain gaps because the network is trained to master the target domain and subsequently the source domain, thus ensuring seamless knowledge transfer. This is attained by introducing the weighting factor $q$ to portray the similarity between the two domains.
@@ 4.4 Algorithm
4.5 Time-complexity Analysis
@@
5 Experiments
5.1 Datasets Specifications
To evaluate the effectiveness of the proposed method, we use a widely used, publicly available CDFSL for HSI classification datasets. We use the Chikusei dataset (CH) yokoya2016airborne for the source domain. Moreover, we use the Indian Pines (IP), Salinas (SA), and Pavia University (PU) datasets for the target domain Li2021DeepCF.
The CH dataset was collected using the Headwall Hyperpec-VNIR-C sensor in Chikusei, Ibaraki, Japan, on July 29, 2014, which comprises 128 bands spanning 340-1018 nm. The image is 2517x2335 and contains 19 categories. More details are provided in table˜1.
The IP dataset was acquired by AVIRIS in 1992 over Northwest Indiana and contains 200 bands spanning 400-2500 nm, with a spatial size of 145x145 and 16 classes. The SA dataset was taken by the AVIRIS sensor in the Salinas Valley, California, USA. The image size is 512x217. The spectral image contains 224 bands, of which 204 bands are useful. The PU dataset was taken by the ROSIS sensor over Pavia, in northern Italy, in 2003. It contains 103 bands after discarding the noise bands, and covers the range 430-860 nm. The image size is 620x340 and has 9 classes. More details about the target datasets are provided in Tables 1-5 Shi2025MultiGaussianPM; Li2024CrossDomainFH.
Table 1: Class names and samples per class for Chikusei dataset
@@ 19 Paved ground 145
Table 2: Class names and samples per class for the Indian Pines dataset
@@ 16 Stone Steel Towers 93
Table 3: Class names and samples per class for the Salinas dataset
@@ 16 Vinyard untrained trellis 1807
Table 4: Class names and samples per class for the Pavia University dataset
@@ 9 Shadows 947
Table 5: Number of training and testing set samples per class in the Houston dataset
@@ 15 Running Track 187 473
5.2 Baseline methods
To evaluate the effectiveness of our method, we comprehensively compare it against the MGPDO Shi2025MultiGaussianPM, CFSL-KT Huang2023HyperspectralIC, CDLA Yu2025FewshotLF, CF-Trans Zhu2025FromIT, CMTL Cheng2023CausalML, GMTN Wang2023GraphMT, DCFSL Li2021DeepCF, CMFSL Xi2022FewShotLW, HFSL Wang2022HeterogeneousFL, CNN-IT Huang2020HyperspectralIC, Gia-CFSL Zhang2022GraphIA, SSRN Zhong2018SpectralSpatialRN, MCNN zheng2020hyperspectral, SPRN zhang2021spectral, FDFSL Qin2024CrossDomainFL, ADAFSL Ye2023AdaptiveDF, RPCL-FSL liu2023category, CASCL Li2024CrossDomainFH, GLGAT-CFSL Ding2024GLGATCFSLGG, and DA-CFSL Qin2025DualalignmentCF.
5.3 Implementation Details
The experiments are conducted on a workstation with an NVIDIA GeForce RTX 4090 GPU and Intel Core i9-14900K. We performed the experiments with 20 episodes (random seeds). We utilize the ViT-B model from HyperSIGMA as the backbone with two ViT branches pre-trained with masked autoencoder (MAE) He2021MaskedAA on the HyperGlobal-450K dataset Wang2024HyperSIGMAHI. Moreover, we use Principal Component Analysis (PCA) to decrease the number of bands to 50 for all datasets as a preprocessing step.
Table 6: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Indian pines dataset (Five labeled samples per class)
@@ 5.3 Implementation Details
SSRN Zhong2018SpectralSpatialRN 99.01 42.91 58.84 81.49 78.63 95.75 99.55 76.91 100 60.51 51.14 53.59 99.29 87.98 75.21 97.71 $65.72\pm 3.77$ $79.22\pm 1.62$ $61.72\pm 3.90$
MCNN zheng2020hyperspectral 93.24 46.14 46.24 76.57 66.39 87.15 97.27 78.14 100 65.32 62.34 53.17 97.14 87.24 75.96 97.58 $66.17\pm 3.21$ $80.32\pm 2.83$ $62.57\pm 3.67$
SPRN zhang2021spectral 98.94 54.76 51.14 83.78 73.53 90.91 99.65 84.13 100 67.95 65.22 49.22 97.58 90.22 80.18 96.51 $67.32\pm 2.08$ $81.28\pm 1.12$ $63.32\pm 2.93$
DCFSL Li2021DeepCF 97.8 45.14 46.88 75.60 67.74 86.50 99.57 76.03 100 57.08 61.45 48.69 95.35 89.48 67.72 96.82 $64.89\pm 3.50$ $75.75\pm 2.44$ $60.36\pm 3.96$
Gia-CFSL Zhang2022GraphIA 93.41 47.98 49.35 76.51 66.65 84.14 99.57 83.76 100 58.23 61.32 45.83 96.45 87.83 70.21 96.02 $65.44\pm 2.23$ $76.08\pm 1.48$ $61.02\pm 2.53$
FDFSL Qin2024CrossDomainFL 96.59 55.62 64.05 97.07 75.17 83.52 99.57 83.11 100 66.10 64.51 59.29 98.70 88.60 80.21 98.18 $71.11\pm 2.19$ $81.27\pm 1.84$ $67.55\pm 2.44$
ADAFSL Ye2023AdaptiveDF 100 51.03 64.10 81.21 79.75 92.65 100 92.30 100 59.54 69.93 59.93 97.20 87.86 71.50 97.73 $71.91\pm 5.15$ $81.54\pm 1.49$ $68.23\pm 3.39$
RPCL-FSL liu2023category 98.54 62.58 57.09 93.23 78.16 90.06 100 90.51 100 73.80 69.56 64.40 98.50 90.55 80.63 97.05 $75.11\pm 1.97$ $84.04\pm 1.10$ $71.94\pm 2.17$
CFSL-KT Huang2023HyperspectralIC 98.57 44.86 72.30 88.45 74.94 96.80 100.0 99.24 100.0 78.82 73.99 61.87 100.0 91.71 83.66 99.77 $76.29\pm 1.26$ $85.36\pm 1.21$ $73.41\pm 1.39$
GLGAT-CFSL Ding2024GLGATCFSLGG 100 78.83 91.93 95.79 96.99 95.03 99.74 98.91 99.67 79.82 80.96 70.03 98.75 97.84 68.93 100 78.5 82.96 79.98
CDLA Yu2025FewshotLF 99.27 68.01 73.56 92.24 80.77 93.23 100.0 95.18 100.0 71.61 70.99 71.43 97.90 92.49 91.39 98.07 $78.94\pm 3.88$ $87.26\pm 1.59$ $76.26\pm 4.23$
DA-CFSL Qin2025DualalignmentCF 99.27 73.23 73.78 91.25 82.34 91.75 100 94.29 100 71.32 70.87 74.78 97.9 90.71 84.78 98.75 $79.26\pm 3.45$ $87.19\pm 1.29$ $76.62\pm 3.73$
CASCL Li2024CrossDomainFH 97.56 73.11 62.85 95.32 77.86 94.86 100 95.32 100 76.25 78.14 63.20 98.33 89.13 86.75 97.78 $79.73\pm 1.48$ $86.34\pm 1.73$ $76.53\pm 1.86$
MGPDO Shi2025MultiGaussianPM 100 71.98 63.56 90.37 83.47 96.37 100 99.93 100 75.32 80.98 59.41 99.33 92.67 91.69 95.45 $81.32\pm 1.62$ $87.53\pm 0.31$ $78.75\pm 1.73$
@@
Table 7: Class-specific classification accuracy (%), OA (%), AA (%), and KC (%) on the Pavia University dataset (Five labeled samples per class)